OpenStudy (anonymous):

Let f(x) = 3x+2 and g(x) = x^2 Part 1 [2 points] Find f(g(x)) Part 2 [2 points] Find g(f(x)) Part 3 [4 points] Use complete sentences to explain any difference in part 1 and part 2.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Mandre

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

f(g(x))=3x^2+2

OpenStudy (anonymous):

How do I solve this?

OpenStudy (phi):

Find f(g(x)) means start with f(x) = 3x+2 replace x with g(x): erase the x, and put in g(x) f(g(x))= 3g(x) + 2 but g(x) is x^2 so replace g(x) with x^2 : f(g(x))= 3x^2 +2

OpenStudy (anonymous):

Okay, does that get simplified?

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (phi):

you can't do anything with 3x^2+2 there are no *like terms*

OpenStudy (anonymous):

Okay, what about part 2?

OpenStudy (phi):

now try g(f(x)) start with g(x) = x^2

OpenStudy (anonymous):

I don't know...

OpenStudy (anonymous):

@phi I don't get it

OpenStudy (phi):

you want g(f(x)) which is short for "start with g(x)" and replace x with f(x)" g(x) = x^2 if you erase the x, and put in f(x) (in other words, put in "f(x)" ) what do you get ?

OpenStudy (anonymous):

(3x+2)^2?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Okay, how do I do part 3? What's different? g(f(x)) is just f(x) squared?

OpenStudy (phi):

if we expand g(f(x))= (3x+2)^2= 9x^2 +6x+4 and compare to f(g(x))= 3x^2+2 they are both quadratics, by g(f(x)) is "bigger" because of the 9x^2 instead of the 3x^2